OpenStudy (ishipdestiel):

Will medal fan and add testimonial A point A is graphed at (0,3). Complete the coordinates of a shape that can be rotated about the y- axis to create a cylinder. Explain

OpenStudy (ishipdestiel):

@rebeccaxhawaii @jfernandes @MilkNCookies @smartnerd1111 @welshfella

OpenStudy (smartnerd1111):

So all the points that satisfy x and y such that x^2 + y^2 = a^2 and -h/2+3 <= z <= h/2 +3 would create a cylinder about the y-axis of radius a and height h.

OpenStudy (ishipdestiel):

So what should I put as my answer and how do I explain it

OpenStudy (smartnerd1111):

that is how you describe a cylinder but i dont know anything after this.i did this before but a while ago

OpenStudy (ishipdestiel):

umm ok thanks